the angular momentum of an electron in 2s orbit is?

Grade:12

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Pankaj

 L = n{h \over 2\pi} = n\hbar
n=2
L=h/∏

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

rajan jha
13 years ago

angular momentum(mvr)=h/2pi(l(l+1))1/2

                                    =h/2pi(0(0+1))1/2             {for 2s orbital l=0}

                                    =0

like my answer?

vikas askiitian expert
13 years ago

 according to bohr's theory ,angular momentam of electron is mvr =nh/2pi

          angular momentam =nh/2pi

  for 2s orbital value of principal quantam number (n) is 2

   therefore angular momentam of electron is 2h/2pi= h/pi....